Hörsaalzuteilung für die Klausur am 14.12.2002

A–F Hörsaalgebäude II, Hörsaal 3

Abker, Melanie Ackermann, Nick Adelt, Henriette Arens, Gregor Böhm, Sebastian Büttner, Christian Badke, Sven Batsch, Bianca Bauer, Kai Baumgartner, Karin Becker, Jana Bensch, Jessica Benthaus, Vanessa Besendorf, Barbara Bialaschik, Kamila Blom, Hendrik Blümel, Florian Blum, Andreas Bolk, Dennis Bolotina, Anna Borchard, Eva Bozkurt, Gabriel Brede, Oliver Brennecke, Peter Bresinsky, Dirk Brinkmann, Britta Brinkmann, Marina Bublitz, Denise Budde, Siegfried Bulla, Matthias Campese, Simon Carl, Stefan Cioica, Petru Adrian Civis, Anna-Katharina Classen, Christopher Dieterich, Janina Dittmer, Benjamin Dittrich, Daniela Domanski, Dariusz Dombrowski, Vitalis Dommers, Andrea Düpmeier, Sara Duscha, Vicki Endt, Dennis Fiebranz, Silke Fischer, Tim Flügel, Thorsten Fuchs, Sabrina Funke, Katharina

G–Sch Audimax

Gajdikova, Magdalena Giese, Vanessa Göckeler, Christian Gröner, Philipp Grob, Andreas Güler, Yeliz Gunaceur, Boutaina Gurris, Marcel Gvoy Sok Ly Hahn, Jutta Hahn, Miriam Hallek, Sandra Harzheim, Louisa Heßdörfer, Jan Hein, Doris Heldt, Daniel Hengesbach, Olaf Hesse, Benedikt Hettermann, Tim Hielen, Gordon Hilsmann, Marie Hörter, Sandra Hövelmann, Karin Hua Zhong Hunkenschroer, Thomas Ibel, Sebastian Idmann, Igor Iffländer, Rebecca Jin, Han Jozwiack, Michael Juchheim, Carsten Jung, Thorsten Jung, Volker Just, Stefanie Kabbash, Moukarram Kang Le Karsten, Alex Keil, Patrick Kemna, Christian Klatt, Jessica Klempan, Johannes Knecht, Melanie König, Moritz Koning, Jennifer Kordel, Matthias Kortmann, Robin Kozakowski, Daniel Krasanskaite, Raimonda Kreysing, Anna Kröger, Julia Krompietz, Zaneta Krüger, Maren Kruse, Alexander Krynen, Bastian Kühbacher, Christian Kunina, Tetyana Kunze, Kathleen Langenscheidt, Dominik Lin Lijun Linnenbrink, Lukas Ljustina, Sretan Lusebrink, Björn Machnik, Matthias Maerus, Marcus Malewski, Sonja Malinowska, Monika Mandapathil, Maneesh Mann, Dominic Mantei, Nadine Marzodko, Barbara Mehanna, Benjamin Messer, Joachim Mitze, Mark-Oliver Müller, Stefanie Müller, Thomas Neveling, Sarah Niehof, Britta Niggemann, Linda Nimmergut, Robert Olah, Edina Ounaceur, Boutaina Özdemir, Ercan Pastewka, Raphaela Philipp, Aylin Piechota, Peter Pielsticker, Judith Plaza, Patric Plaza, Robert Podzielny, Patrick Preger, Sven Proß, Niklas Pütter, Daniel Qu, Qingnian Quade, Peter Raffler, Annika Ralser, Tobias Rehling, Dirk Reimann, Sarah Reiners, Michael Reinhard, Sven Rempe, Jasmin Röglin, Heiko Roth, Alfred Rozgic, Marco Sawlanski, Stefan Schädel, Lara Schartner, Inna Scheffer, Marcus Scheimann, Swetlana Schewe, Katharina Schiebold, Lars Schlinkmann, Verena Schmelter, Mailin Schmidt, Lars Schmidt, Yvonne Schmitz, Ines Schmitz, Lisa Maria Schönfelder, Frank Schowe, Benjamin Schröder, Nadine Schubert, Melanie Schütze, Mario Schumacher, Anne Schwanengel, Bernadette Schwietz, Roland

Se–Z Hörsaalgebäude II, Hörsaal 5

Sechi, Stefania Sender, Julia Sendes, Juliane Sewcz, Melanie Sibin, Nelli Simon, Matthias Singh, Sanjay Spieler, Thomas Spira, Gerrit Sprengel, Sabrina Springer, Tobias Steinwes, Nina Stotko, Dominik Stracke, Dennis Sun Ning Thimm, Matthias Timm, Thomas Tombrinck, Jennifer Twieling, Gert Uesseler, Peter Marco Vogelsang, Gesa Vogt, Nikolas Volkers, Gisela Voß, Stefan Wagner, Sintje Walkowiak, Brigitte Walter, Katrin Wegener, Olaf Wegener, Yvonne Wegner, Sebastian Weidner, Sandra Weigand, Helene Weinberger, Regina Wember, Rebecca Weßel, Stephanie Weyland, Dennis Wieczorek, Jakob Wieschnewski, Alexander Wiesollek, Marc Windt, Anna Wohlgemuth, Andrea Wolf, Jan Woroniuk, Katharin Xing Wei Zacher, Sabine Zaytseva, Eugeniya Zhang, Yan Zhong, Hua Zwiener, Marc